OpenStudy (anonymous):

(1 - sin x)(1 + sin x) = cos^2 x verify the identity

OpenStudy (anonymous):

\[(1 - \sin x)/(1 + \sin x) = cosx/( 1+sinx)\]

OpenStudy (anonymous):

\[(1-b)(a+b)=1-b^2\]

OpenStudy (anonymous):

ca you explain that to me?

OpenStudy (anonymous):

second "identity" you wrote is wrong

OpenStudy (anonymous):

i re wrote it as the first answer thing:)

OpenStudy (anonymous):

this \((1 - \sin( x))(1 + \sin (x)) = \cos^2 (x)\) is what you are trying to verify, right?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

oh sorry

OpenStudy (anonymous):

(1-sinx)/cosx=cosx/(1+sinx)

OpenStudy (anonymous):

Remember that \[\cos^{2}(x) + \sin^{2}(x) = 1\]

OpenStudy (anonymous):

thats it @satellite73

OpenStudy (anonymous):

i dont understand that one

OpenStudy (anonymous):

@mebs

OpenStudy (anonymous):

can you please explain it to me ?